# Impurities in a trapped 1D Bose gas of arbitrary interaction strength: localization-delocalization transition and absence of self-localization

### v1

@@ II.2 DMRG simulation of quantum impurity in a 1D Bose gas
Here the terms containing $\hat{a}^{\dagger}_{i}\hat{a}_{i+1}$ ($\hat{b}^{\dagger}_{i}\hat{b}_{i+1}$) describe the hopping between lattice sites for the bosons (impurity) with hopping amplitude $J$ ($J_{\mathrm{I}}$.) The diagonal terms $\sim\hat{a}^{\dagger}_{i}\hat{a}_{i}$ (or $\hat{b}^{\dagger}_{i}\hat{b}_{i}$) describe a local potential at site $i$ for the bosons (impurity). The strength of this potential is also given by the hopping amplitude $J$ ($J_{\mathrm{I}}$). It doesn’t effect the dynamics of the system but needs to be taken into account when calculating the ground state energy. The terms proportional to $\hat{a}^{\dagger}_{i}\hat{a}^{\dagger}_{i}\hat{a}_{i}\hat{a}_{i}$ ($\hat{b}^{\dagger}_{i}\hat{a}^{\dagger}_{i}\hat{a}_{i}\hat{b}_{i}$) relate to the interaction between the bosons (bosons and impurity). The strength of this interaction is given by $U$ ($U_{\mathrm{I}}$). A graphical illustration of the terms of Hamiltonian eq. (6) is given in Fig. 2a).
Abbildung 2: a) Illustration of terms in Hamiltonian eq. (6). The blue circles depict bosons which interact with strength $\frac{U}{2}$ and can hop with amplitude $J$. The impurity is shown as a red circle. It interacts with the bosons with strength $U_{\mathrm{I}}$ and hops between lattice sites with amplitude $J_{\mathrm{I}}$. b) Comparison between the sinusoidal dispersion relation of particles in a lattice to that of free particles, with the maximal energy to which both remain comparable.
@@
Abbildung 3: Comparison of the saturated polaron energy ($g_{\text{IB}}=\infty$) for an immobile impurity ($M=\infty$) in a Bose gas with $N=100$ particles from DMRG and diffusion quantum Monte-Carlo results (DMC) from [9].
@@
In order to benchmark our DMRG code we calculated the saturated polaron energy ($g_{\text{IB}}=\infty$) of an infinite-mass polaron $(M=\infty)$ localized in the center of a 1D quasi condensate with $N=100$ bosons for varying strength of the boson-boson interaction, quantified by $\gamma$, and compared the results with those from diffusion quantum Monte Carlo (DMC) simulations taken from [9], see Fig. 3. The DMC simulations were done for periodic boundary conditions, while the DMRG data are for open boundary conditions, which explains the small difference for small value of $\gamma$. Open boundary conditions cause half dark solitions to form on each system edge for $\gamma\ll 1$ and Friedel oscillations for $\gamma\gg 1$ which increase the ground state energy of the system compared to periodic boundary conditions. This was mostly negated here by choosing a system size which is large compared to the characteristic length scale of the system, the healing length $\xi=1/\sqrt{2gn_{0}m}$ for $\gamma\ll 1$ or the wavelength of Friedel oscillations $2k_{F}=2\pi n_{0}$ for $\gamma\gg 1$. Apart for very small $\gamma$ values one recognizes excellent agreement.
II.3 Decoupling mean-field (DMF) ansatz
$\quad+g|\phi(x,t)|^{2}+V_{B}(x)\biggr{)}\phi(x,t),$ (12)
III Localization-Delocalization transition in a box potential
We now consider a box potential for bosons and impurity $V_{B}(x)=V_{I}(x)$ for which we expect a transition between a delocalized phase, where the impurity is spread out over the whole system, and a localized one, where the impurity is localized at one of the two edges of the system. This is because the energy of an impurity localized at the edge increases as $E_{\textrm{loc}}\sim\sqrt{g_{\text{IB}}}$, while the polaron energy in the bulk scales as $E_{p}\sim g_{\text{IB}}$. Thus for $g_{\text{IB}}\to 0$ no bound state exists which only emerges above a critical value of $g_{\text{IB}}$.
In [30, 31] it has been argued based on a DMF ansatz, eq. (12), that also in a translational invariant, homogeneous system the impurity wavefunction should self-trap in a co-localized distortion of the BEC. The reality of such self-trapped polarons has been subject of discussions [32]. Due to the broken translational invariance in the box potential considered here, the self-trapping would manifest itself in a probability density of the impurity centered in the middle of the trap with a width independent on the trap size $L$.
III.1 Localization-delocalization transition in a box potential
We performed DMRG simulations of the ground state of a mobile impurity in a 1D Bose gas in the box potential for different mass ratios $M/m$ and impurity-boson interaction strengths $g_{\text{IB}}$. Results for weak boson-boson interactions, $\gamma=0.4$, and for the Tonks limit $\gamma=\infty$ are shown in Fig. 4 and Fig. 5, respectively. One recognizes in both cases a sharp transition between a phase where the impurity is delocalized over the trap to a phase where it is localized at one of the two edges. (Note that the true ground state is a superposition of localized states at both edges. Since the energy difference between the symmetric and antisymmetric superpositions vanishes exponentially with increasing system size, the DMRG algorithm converges to a solution on one edge only.) The transition point depends on the boson-boson interaction strength. In the delocalized phase the probability density of the impurity is spread out over the whole system and apart from some corrections at the edges it corresponds to the ground state of the box potential $\sim\cos^{2}(x\pi/L)$, with $L$ being the length of the box. In the localized phase, on the other hand, the width of the probability distribution becomes independent on system size for large $L$.
Abbildung 4: a) Localization-delocalization transition of impurity in a box potential in a weakly interacting Bose gas with Tonks parameter $\gamma=0.4$ and $N=40$ particles. Color code describes average position $\langle\hat{X}\rangle$ of impurity. Solid line is the prediction from eq. (15) with $l_{0}=\xi$. The dashed line gives the prediction with $l_{0}$ from a variational ansatz. b) and c) density distribution corresponding to the two points in Figure a).
$E_{p}\approx g_{\text{IB}}n_{0}$ (13)
and thus scales linearly in $g_{\text{IB}}$.
For a weakly interacting Bose gas (small $\gamma$) the characteristic length scale $l_{0}=\eta\xi$ of the harmonic confinement is the healing length $\xi$ of the Bose condensate. The factor $\eta\simeq{\cal O}(1)$ accounts for the backaction of the trapped impurity on the density of the condensate near the potential edge, which is relevant in the small-$\gamma$ regime, due to the large compressibility of the weakly interacting bose gas. The impurity thus experiences an effective potential for small $x$:
$\phi(x)=\sqrt{n_{0}}\tanh\left(\alpha x\right).$ (16)
describing the coherent amplitude of the condensate at the edge. $|\phi_{I}\rangle$ is then the ground state of the impurity in the resulting effective Pöschl-Teller potential, which can be calculated analytically [33]. Minimizing the total energy gives an implicit equation for $\eta=1/(\sqrt{2}\xi\alpha)$ as function of the mass ratio $M/m$, the Tonks parameter $\gamma$ and the normalized Boson-impurity interaction strength $\gamma_{\textrm{IB}}=g_{\text{IB}}m/n_{0}$
Abbildung 5: a) Localization-delocalization transition of an impurity in a box potential in a Tonks gas ($\gamma=\infty$, $N=40$). Solid line is prediction from eq. (19). b) and c) density distribution corresponding to the two points in Figure a).
$n(x)\approx\frac{2k_{F}}{\pi}\left(1-\frac{2\sin(2k_{F}x)}{2k_{F}x}\right),$ (18)
with $x$ being the distance from the edge and $k_{F}=\pi n_{0}/2$ is the Fermi momentum. Thus the characteristic length scale $l_{0}$ is here $(\sqrt{6}/\pi)n_{0}^{-1}$ which gives the following estimate for the critical value of the localization-delocalization transition:
III.2 Absence of self-localization in a weakly interacting Bose gas
Abbildung 6: Comparison between average impurity position $\expectationvalue{\hat{X}}$ calculated from DMRG and decoupling mean field (DMF) calculations for a Bose gas with $\gamma=0.4$ and $N=40$. a) Colorplot of $\expectationvalue{\hat{X}}$ as function of mass ratios ${M}/{m}$ and bose-impurity interaction $g_{\text{IB}}$ obtained by DMRG. Red lines show cuts which are plotted in c). b) The same obtained from decoupling mean field simulations (DMF). c) Cuts in a) for two mass ratios, solid and dashed line correspond to each other.
Abbildung 7: Comparison between the variance of the impurity $\langle\Delta\hat{X}\rangle$ calculated from DMRG and decoupling mean field calculations for a Bose gas with $\gamma=0.4$ and $N=40$. a) Colorplot of variance from DMRG simulations as function of mass ratios ${M}/{m}$ and boson-impurity interaction $g_{\text{IB}}$. Red lines show cuts which are plotted in c). b) The same obtained from decoupling mean field simulations (DMF). c) Cuts in a) for two mass ratios, solid and dashed line correspond to each other. The grey dashed-dotted line shows the spatial width of the impurity expected for $g_{\text{IB}}=0$, which is given by the ground-state density distribution of a particle in a box $\cos^{2}{\left(x\pi/L\right)}$. The black dotted line shows the self-localization length $\lambda$ predicted in DMF for $\frac{M}{m}=1.11$.
Abbildung 8: Left: Density distribution of Bose gas and impurity for different $g_{\text{IB}}$ (written in the lower right) from DMRG simulations. Right: The same from mean-field calculations compared to the analytic prediction for a self-trapped polaron, eq. (20). Parameters of the bose gas are $\gamma=0.4$, $M/m=1$ and $N=40$.
[…]
The self-localisation is an artifact of the decoupling approximation and is not present in the full quantum approach. This is because boson-impurity correlations, which are neglected in DMF, suppress the self-localization [32].
IV Heavy polaron in strongly interacting bose gas
A characteristic quantity of the polaron is the energy $E_{\mathrm{P}}$ needed to immerse an impurity into the Bose gas $E_{\mathrm{P}}=E(g_{\text{IB}}\neq 0)-E(g_{\text{IB}}=0)$. Here $E(g_{\text{IB}}=0)$ is the ground state energy of the system without Bose-impurity-interaction and $E(g_{\text{IB}}\neq 0)$ with Bose-impurity-interaction.
In the following we will employ DMRG simulations to calculate the polaron energy in a strongly interactig Bose gas.
IV.1 Polaron energy and density profile
For an impurity whose mass is much larger then that of the bosons ($\frac{M}{m}\gg 1$) the kinetic energy of the impurity can be neglected. The resulting effective Hamiltonian then reduces to a interacting boson problem in an external $\delta$ potential
Figures 9 and 10 show the polaron energy in a weakly ($\gamma=0.4$) and a strongly ($\gamma=20)$ interacting, trapped 1D Bose gas, calculated by DMRG simulations. For $\gamma=0.4$ the analytic prediction from eq. (21) is plotted for comparison, showing very good agreement. Also shown are plots for the boson density for two different values of $g_{\text{IB}}$. In the first case the characteristic size of the condensate depletion around the fixed position of the impurity is given by the healing length, in the second case by the wavelength of the Friedel oscillations of the (almost) Tonks gas. One recognizes that in both cases the polaron energy saturates for increasing $g_{\text{IB}}$, which happens when the density of the Bose gas at the position of the impurity approaches zero.
Abbildung 9: Polaron in weakly interacting 1D Bose gas with $\gamma=0.1$, $M=\infty$ and $N=80$: a) polaron energy as function of impurity-boson interaction calculated from DMRG simulations and Meanfield approximations in the LLP frame [11, 12], b), c) wavefunction for two different values of $g_{\text{IB}}$ (written in the lower right) in a box and impurity fixed at $x=0$.
Abbildung 10: Polaron in strongly interacting 1D Bose gas with $\gamma=20$ and $N=80$ particles in a finite box and $M=\infty$: a) polaron energy as function of impurity-boson interaction, b), c) wavefunction for two different values of $g_{\text{IB}}$ (written in the lower right) in a box potential and impurity fixed at $x=0$.
IV.2 Analytic approximation of the polaron energy for $\gamma\gg 1$
$\hat{H}_{F}$ $=\int\text{d}x\,\hat{\psi}^{\dagger}(x)\left(-\frac{\partial^{2}_{x}}{2m}+g_{\mathrm{IB}}\delta\left(x\right)\right)\hat{\psi}(x)+$ (23)
$-\frac{g_{\mathrm{F}}}{2}\int\!\!\!\int\text{d}x\,\text{d}y\,\,\hat{\psi}^{\dagger}(x)\hat{\psi}^{\dagger}(y)\overset{\leftarrow}{\frac{\partial}{\partial z}}\delta(z)\overset{\rightarrow}{\frac{\partial}{\partial z}}\hat{\psi}(y)\hat{\psi}(x),$
$g_{\mathrm{F}}=-\frac{4}{g}.$ (24)
The boson-fermion mapping provides an elegant way to find approximate analytic expressions for the polaron energy in a strongly interacting Bose gas with $\gamma\gg 1$ perturbatively in $|g_{\mathrm{F}}|\sim 1/\gamma$. To this end we first have to calculate the single particle solutions of free fermions with a infinitely heavy impurity at $x=0$ for open boundary conditions (OBCs). This system is described by the Hamiltonian
$\hat{H}_{F}^{0}=\int_{-L/2}^{L/2}\!\!\!\!\text{d}x\,\,\hat{\psi}^{\dagger}(x)\left(-\frac{\partial^{2}_{x}}{2m}+g_{\text{IB}}\delta\left(x\right)\right)\hat{\psi}(x).$ (25)
This means we need to solve the problem of non-interacting fermions trapped in a infinitely deep box potential (OBC’s) with a delta potential in the middle.
The single particle solutions are given by
$\varphi_{l}(x)=a_{l}\mathrm{sgn}(x)^{l}\sin\left(k_{l}\left(|x|-\frac{L}{2}\right)\right)$ (26)
where $l$ is an integer and the wave numbers $k_{l}$ have to be numerically calculated from the boundary condition at the delta potential
$E=$ $\sum_{l=1}^{N}\frac{k_{l}^{2}}{2m}.$ (29)
$\Delta E_{1}=\bra{\psi}\hat{H}_{1}\ket{\psi}$ (30)
Fig. 11 shows a comparison of the polaron energies as function of $g_{\text{IB}}$ from first order perturbation in $1/\gamma$ with numerical DMRG results. One recognizes rather good agreement even down to $\gamma=20$.
[…]
For a weakly interacting Bose gas $(\gamma\ll 1)$ a mean-field ansatz in the LLP frame can be used to obtain semi-analytic expressions of the interaction potential at short distances, which agree very well with quantum Monte-Carlo simulations [13]. The mean field approach does not describe the Casimir-like contributions arizing from the exchange of virtual phonons [44, 45, 46], which give however only important corrections in the already small tails of the interaction potential.
In this section we investigate the interaction between Bose polarons in a strongly-interacting 1D quasi condensate numerically by DMRG simulations. Since DMRG is not well suited for periodic boundary conditions, we again assume a confinement of the Bose gas to a box potential. As long as the distance of the impurities from the edges of the box is much larger than the healing length or the Fermi wavelength, boundary effects can be neglected. For the same reason our simulations, although in principle suitable, do not allow to accurately extract the far tails of the interaction potential, dominated by virtual phonon exchange. As discussed in the Appendix, the DMRG simulations reproduce the semi-analytical results from Ref.[13] obtained in mean-field theory for small values of $\gamma$.
$V(r)=E_{lr}(r)-E_{l}(r)-E_{r}(r)+E_{0}.$ (32)
Abbildung 12: Left: Interaction potential of heavy impurities in a bose gas with $\gamma=4$ and $N=100$ particles and bose-impurity interaction strength $g_{\text{IB}}m/n_{0}=1.08$. Right: Density of Bose gas at the two separations indicated in left picture as crosses in the corresponding color.
Abbildung 13: Left: Interaction potential of heavy impurities in the Tonks-gas limit $\gamma=\infty$, $N=80$ and $g_{\textrm{IB}}m/n_{0}=1.58$. Insert illustrates origin of oscillatory potential. Also shown is the analytic prediction (black line) of the interaction potential obtained from mapping to free fermions in PBC. Right: Density of Bose gas at the two separations indicated in left picture as crosses in the corresponding color.
For large values of $\gamma$ on the other hand oscillations appear in the polaron potential, as can be seen in Fig. 13. The frequency of these oscillations is that of the Friedel oscillations caused by the polarons themselves. As shown in Fig. 13b.), once a single impurity is introduced into the system it causes density oscillation and the second impurity has to displace a smaller or a larger amount of bosons and therefore needs less or more energy to form a polaron. The resulting energy modulations reflect themselves in the polaron potential.
$V(r)=E_{\mathrm{lr}}-2E_{1}+E_{0}.$ (34)
As can be seen in Fig. 13 the bipolaron potential as well as the density distribution of the Tonks gas obtained in this way compare quite well with the DMRG simulations.
VI Summary
In the present paper we studied the ground state of a mobile impurity in a one-dimensional Bose gas in a box potential for arbitrary impurity-boson and boson-boson interactions. To fully account for quantum effects, which are particularly important in the limit of large Tonks parameters $\gamma$, we performed numerical simulations of a discretized effective lattice model using the density-matrix renormalization group.
Mean-field approaches to the Bose polaron have been shown to provide accurate descriptions for heavy impurities ($M/m\geq 3$) and weak boson-boson interactions $(\gamma\leq 1)$, if the back-action of the impurity to the condensate is properly taken into account. We here showed that it is important to also keep impurity-boson correlations in the mean-field description, in agreement with recent findings in [32]. In translationally invariant systems or for harmonic trapping of both impurity and bosons this can be achieved by a Lee-Low-Pines transformation of the original Hamiltonian, reducing the interaction problem to one of interating bosons in a delta-Potential, before applying a mean-field approximation. On the other hand, a commonly used mean-field approach, where condensate and impurity are described by a factorized impurity-boson c-number wavefunction, leading to a coupled Gross-Pitaevski – Schrödinger equation, disregards such correlations. As a consequence such an approach predicts the existence of a self-trapped polaron for arbitrarily small impurity-boson couplings in a homogeneous 1D gas, where the impurity is localized in a distortion of the condensate created by the impurity. A comparison of the predictions of such a decoupling mean-field (DMF) theory with exact DMRG results shows that the self-trapped solution is an artifact of the DMF approach.
As DMRG simulations allow us to consider Bose polarons in a strongly-interacting Bose gas, we furthermore calculated the polaron energy in this regime and derived analytical approximations for large but finite Tonks parameters $\gamma\gg 1$ by using the mapping to weakly interacting fermions. Finally we numerically calculated the short-distance interaction potential between two impurities in Born-Oppenheimer approximation for arbitrarily strong boson-boson interactions. For small Tonks parameter $\gamma\leq 1$ we verified the results of [13] were a linear short-distance behaviour was predicted. In the strong interaction limit $\gamma\gg 1$ we found oscillatory modulations in the potential, which can be interpreted as the effect of Friedel-type oscillations in the condensate density caused by one impurity affecting the polaron formation of the second impurity.
Acknowledgement
The authors gratefully acknowledge financial support by the DFG through SFB/TR 185, Project No.277625399.
Appendix
VI.1 Discretization of continuous gas: Higher order representation of the second order derivatives
Let us comment on some techniques to improve the numerical simulations. It turns out to be beneficial for the convergence of the DMRG algorithm to use a higher order representation of the second order derivatives. So

### v2

@@ VI.1 DMRG simulation of quantum impurity in a 1D Bose gas
Here the terms containing $\hat{a}^{\dagger}_{i}\hat{a}_{i+1}$ ($\hat{b}^{\dagger}_{i}\hat{b}_{i+1}$) describe the hopping between lattice sites for the bosons (impurity) with hopping amplitude $J$ ($J_{\mathrm{I}}$). The diagonal terms $\sim\hat{a}^{\dagger}_{i}\hat{a}_{i}$ (or $\hat{b}^{\dagger}_{i}\hat{b}_{i}$) describe a local potential at site $i$ for the bosons (impurity). The strength of this potential is also given by the hopping amplitude $J$ ($J_{\mathrm{I}}$). It doesn’t effect the dynamics of the system but needs to be taken into account when calculating the ground state energy. The terms proportional to $\hat{a}^{\dagger}_{i}\hat{a}^{\dagger}_{i}\hat{a}_{i}\hat{a}_{i}$ ($\hat{b}^{\dagger}_{i}\hat{a}^{\dagger}_{i}\hat{a}_{i}\hat{b}_{i}$) relate to the interaction between the bosons (bosons and impurity). The strength of this interaction is given by $U$ ($U_{\mathrm{I}}$). A graphical illustration of the terms of Hamiltonian eq. (32) is given in Fig. 12a).
Abbildung 12: a) Illustration of terms in Hamiltonian eq. (32). The blue circles depict bosons which interact with strength $\frac{U}{2}$ and can hop with amplitude $J$. The impurity is shown as a red circle. It interacts with the bosons with strength $U_{\mathrm{I}}$ and hops between lattice sites with amplitude $J_{\mathrm{I}}$. b) Comparison between the sinusoidal dispersion relation of particles in a lattice to that of free particles, with the maximal energy to which both remain comparable.
@@
In addition to the fact that the discrete system needs to be a faithful representation of the continues one, one has also to keep the compression of the DMRG in mind. We utilise the Julia language library ITensors [27] for the implementation of the DMRG algorithm and set the bond dimension compression error to $10^{-6}$. This is sufficiently small such that the state derived from the DMRG is a valid representation of the uncompressed state, but large enough that the system is still computable on a normal PC.
@@
Abbildung 13: Comparison of the saturated polaron energy ($g_{\text{IB}}=\infty$) for an immobile impurity ($M=\infty$) in a Bose gas with $N=100$ particles from DMRG for OBC and diffusion quantum Monte-Carlo results (DMC) for PBC from [9].
@@
In order to benchmark our DMRG code we calculated the saturated polaron energy ($g_{\text{IB}}=\infty$) of an infinite-mass polaron $(M=\infty)$ localized in the center of a 1D quasi condensate with $N=100$ bosons for varying strength of the boson-boson interaction, quantified by $\gamma$, and compared the results with those from diffusion quantum Monte Carlo (DMC) simulations taken from [9], see Fig. 13. The DMC simulations were done for periodic boundary conditions, while the DMRG data are for open boundary conditions, which explains the small difference for small value of $\gamma$. Open boundary conditions cause half dark solitions to form on each system edge for $\gamma\ll 1$ and Friedel oscillations for $\gamma\gg 1$ which increase the ground state energy of the system compared to periodic boundary conditions. This was mostly negated here by choosing a system size which is large compared to the characteristic length scale of the system, the healing length $\xi=1/\sqrt{2gn_{0}m}$ for $\gamma\ll 1$ or the wavelength of Friedel oscillations $2k_{F}=2\pi n_{0}$ for $\gamma\gg 1$. Apart for very small $\gamma$ values one recognizes excellent agreement. For small $\gamma$ the polaron energies obtained with open boundary conditions (OBCs) as done in our DMR simulations are slightly larger than those obtained with periodic boundary conditions (PBCs) used in DMC as the spatial extend of the polaron, which is on the order of the healing length, becomes comparable to the system size:
$\frac{L}{\xi}=\sqrt{2\gamma}N.$ (38)
Let us finally comment on some techniques to improve the numerical simulations. It turns out to be beneficial for the convergence of the DMRG algorithm to use a higher order representation of the second order derivatives. So
@@ VI.2 Comparison of polaron-polaron interaction potential to mean-field result
  • Recati et al. [2005] A. Recati, J.-N. Fuchs, C. S. Peca, and W. Zwerger, Casimir forces between defects in one-dimensional quantum liquids, Physical Review A—Atomic, Molecular, and Optical Physics 72, 023616 (2005).
  • Fuchs et al. [2007] J. Fuchs, A. Recati, and W. Zwerger, Oscillating casimir force between impurities in one-dimensional fermi liquids, Physical Review A—Atomic, Molecular, and Optical Physics 75, 043615 (2007).
@@
  • Petković et al. [2020] A. Petković, B. Reichert, and Z. Ristivojevic, Density profile of a semi-infinite one-dimensional bose gas and bound states of the impurity, Phys. Rev. Res. 2, 043104 (2020).
  • Parisi and Giorgini [2017] L. Parisi and S. Giorgini, Quantum monte carlo study of the bose-polaron problem in a one-dimensional gas with contact interactions, Phys. Rev. A 95, 023619 (2017).
@@
  • Schecter and Kamenev [2014] M. Schecter and A. Kamenev, Phonon-mediated casimir interaction between mobile impurities in one-dimensional quantum liquids, Physical review letters 112, 155301 (2014).
  • Huber et al. [2019] D. Huber, H.-W. Hammer, and A. Volosniev, In-medium bound states of two bosonic impurities in a one-dimensional fermi gas, Physical Review Research 1, 033177 (2019).
  • Volosniev et al. [2015] A. G. Volosniev, H. W. Hammer, and N. T. Zinner, Real-time dynamics of an impurity in an ideal Bose gas in a trap, Phys. Rev. A - At. Mol. Opt. Phys. 92, 023623 (2015).
  • Dehkharghani et al. [2015] A. S. Dehkharghani, A. Volosniev, and N. Zinner, Quantum impurity in a one-dimensional trapped bose gas, Physical Review A 92, 031601 (2015).
  • Gómez-Lozada et al. [2024] F. Gómez-Lozada, H. Hiyane, T. Busch, and T. Fogarty, Bose-fermi $n$-polaron state emergence from correlation-mediated blocking of phase separation, arXiv preprint arXiv:2409.13785  (2024).